Circular No. 7026

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 T1 (LINEAR) One of 17 fast-moving asteroidal objects reported by the Lincoln Lab-

oratory Near Earth Asteroid Research Team on Oct. 2 appears to be a long-period comet, according both to the retrograde nature of its orbit and remarks by Y. Ikari (Moriyama, Japan; relayed by S. Nakano) concerning the object’s appearance, when he observed it on Oct. 3 and 8 in response to the request in The NEO Confirmation Page. From the more complete information given on MPEC 1998-T15, the discovery position, orbital elements and ephemeris are extracted below:

1998 UT  α_{2000}  δ_{2000}  m_{1}
Oct. 2.28667  1^{h}54^{m}23.^{s}56  +23^{o}12′36′′.5  17.6
T = 1999 June 25.625 TT  ω = 225.^{o}855
Ω = 153.122 }2000.0
q = 1.47639 AU  i = 170.188
1998TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 4 1^{h}51.^{m}00 +22^{o}57.^{′}4 2.701 3.629 154.^{o}4 6.^{o}8 15.8
14 1 29.69 +21 09.2 2.553 3.531 166.4  3.8 15.5
24 1 06.69 +18 49.6 2.456 3.432 166.7  3.8 15.3
Nov. 3 0 43.65 +16 06.1 2.413 3.332 153.9  7.5 15.1
13 0 22.17 +13 11.8 2.420 3.233 138.9 11.6 15.0
23 0 03.47 +10 21.6 2.472 3.133 123.9 15.2 14.9

VARIABLE STAR IN HERCULES E. Halderson, T. Shefler, M. Modjaz, J. Y. King, M. Papenkova, W. D.

Li, R. R. Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of a (probably cataclysmic) variable star in our own galaxy during the course of the Lick Observatory Supernova Search with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). The variable star was found and confirmed on unfiltered images taken on Oct. 5.2 (mag ∼ 18.0) and 6.2 UT (mag ∼ 17.8). The object is located at α = 18^{h}14^{m}14^{s}.31, δ = +30^{o}43′37′′.0 (equinox 2000.0). A KAIT image of the same field obtained on Sept. 19.2 shows nothing at the position of the variable star (limiting mag ∼ 19.5), nor does the Digitized Sky Survey image of the field show anything at this position. A finding chart may be found

at http://astron.berkeley.edu/~bait/cv2.gif.

1998 October 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green